Problem 2: Let z = a + jb be a complex number. (In this course, we use j instead of the more common notation i for the imagin2
Answer #1

Solution: z = a + jb = Vo? +62 (22+52 + 1 Ja2 +62) .... (i)
Let 79+zA_ = ours 2? + -= osoo
So, equation (i) can be written as,

2= Va? +62 (coso + jsino) = Va? + b2e1, which is of the form q=relo

So, the magnitude of z = r = Va? +62 and phase of is  o = tan-19